A certain mass of gas has a volume of 1 litre at a temprature of 3degrees celsius and a pressure of 840mm of Hg.What is the volume occupied by it at a pressure of 600mm of Hg at 27degrees celsius?

Grade:11

1 Answers

Nirmal Singh.
10 years ago
P1*V1/T1 = P2*V2/T2
(h1*rho*g)*(1) / (273+3) = (h2*rho*g)*(V2) / (273+27)
840*1 / 276 = 600*V2 / 300
V2 = 1.52 liter
Regards,
Nirmasl Singh
Askiitians faculty